Activity 1 Find the critical value of the following





Lone-tailed test, or known, a= 0.05; n = 34,





2 two-tailed iest, a unknown, a = 0.01, n=23





3 night-tailed test, a unknown; a = 0.01, n = 15





4 one-tailed test, a known, a = 0.025; n = 37





5 left-tled test, a known, a= 0.05, n = 36





1
Expert's answer
2022-05-23T23:14:57-0400

1. Based on the information provided, the significance level is "\alpha = 0.05,"

the critical value for a left-tailed test is "z_c = -1.6449," the critical value for a right-tailed test is "z_c = 1.6449."


2. Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=22" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.624494."


3. Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=14" degrees of freedom, and the critical value for a right-tailed test is "t_c = 2.818756."


4. Based on the information provided, the significance level is "\alpha = 0.025,"

the critical value for a left-tailed test is "z_c = -1.96," the critical value for a right-tailed test is "z_c = 1.96."


5. Based on the information provided, the significance level is "\alpha = 0.05,"

the critical value for a left-tailed test is "z_c = -1.6449."